A man is 6 times as old as his son. 20 years later, the man will be twice as old as his son. Find the age of the man when his son was born.
Answer by Alan3354(69443)

A man is 6 times as old as his son. 20 years later, the man will be twice as old as his son. Find the age of the man when his son was born.
--------------
m = 6s
m+20 = 2*(s+20)
----------------
6s+20 = 2s + 40
s = 5
m = 30
-------
m - s = 25